Grade 11Physical Chemistry

The mass of caco3 is produced when carbon dioxide is bubbled through 500 ml of 0.5M ca(oh)2 will be

Nanesh
9 Years agoGrade 11

1 Answer

Anjana Sirothiya
9 Years ago
Ca(Oh)2 0.5=n/0.5n=0.25Mass=0.25*74=18.5 -------(1)Caco3Wt./100 ------(2)Equating 1and2We get 1850g